OpenStudy (anonymous):

What is the range of the function y = x2?

OpenStudy (anonymous):

you mean y=x^2 or y=2*x?

OpenStudy (anonymous):

if y=x^2 , Df=R and f(Df)=[0,+inf) if y=2*x , f:R->R